OpenStudy (wiggler):

How do I derive 2y^4 - x^4 = 3xy + 4? I get that I need to use implicit differentiation, but keep getting incorrect negative/positive values.

OpenStudy (anonymous):

Wait is that your function and you have to find the derivative of it?

OpenStudy (anonymous):

\[\huge f(x) = 2y ^{4}-x ^{4}=3xy+4\]

OpenStudy (wiggler):

Yeah... it could be rewritten y = (2y^4 - x^4 - 4)/3x

OpenStudy (anonymous):

ok just to make sure again, it's exactly what I wrote right?

OpenStudy (wiggler):

I think so.

OpenStudy (wiggler):

I trying to find dy/dx.

OpenStudy (anonymous):

\[\huge \frac{ d }{ dx }(-x ^{4}+2^{4})=\frac{ d }{ dx }(4+3xy)\] \[\huge 2\frac{ d }{ dx }(y^4)-\frac{ d }{ dx }(x^4) = \frac{ d }{ dx }(4+3xy)\]

OpenStudy (anonymous):

Do the left side than do the right, do you see where this is going?

OpenStudy (anonymous):

The first line should say 2y^4

OpenStudy (wiggler):

Yep, I get \[8y^3 \frac{ dy }{ dx } -4x^3 = 3y + 3x \frac{ dy }{ dx }\]

OpenStudy (wiggler):

Rearranging that gives \[\frac{ dy }{ dx } = \frac{ 3y = 4x^3 }{ 8y^3- 3x }\]

OpenStudy (wiggler):

Opps, on top should be 3y + 4x^3

OpenStudy (wiggler):

But it should be \[- \frac{ 3y-4x^3 }{ 3x-8y^3 }\]

OpenStudy (anonymous):

Alright, sorry about that, I was helping someone else :), ok lets see where you went wrong.

OpenStudy (anonymous):

3y+3x*dy/dx how did you get that? (right side)

OpenStudy (wiggler):

Product rule. ie (3x)'y + 3x(y)' Gives 3y + 3x dy/dx I may have done that completely wrong.

OpenStudy (anonymous):

Alright I'm just going to do the right side for you step by step.

OpenStudy (anonymous):

\[\frac{ d }{ dx }(4)+3\frac{ d }{ dx }(xy)\] \[3(\frac{ d }{ dx }(xy))+ 0\] \[3x \frac{ d }{ dx }(y)+\frac{ d }{ dx }(x)(y)\] \[3(\frac{ dy }{ dx }(x)+(\frac{ d }{ dx }(x))y)\] \[3y+3x \frac{ dy }{ dx }\] Ok so this is where you got to I see and had trouble.

OpenStudy (wiggler):

Yep, that's exactly what I did.

OpenStudy (anonymous):

\[\huge -4x^3+8y^3\frac{ dy }{ dx }=3y+3x \frac{ dy }{ dx }\] \[\huge -4x^3-3x \frac{ dy }{ dx }+8y^3\frac{ dy }{ dx }=3y\]

OpenStudy (anonymous):

The whole point of these questions is to isolate dy/dx, can you finish it off?

OpenStudy (wiggler):

I still get what I was getting before: \[8y^3\frac{ dy }{ dx } - 3x \frac{ dy }{ dx } = 3y+4x^3\] \[(8y^3-3x)\frac{ dy }{ dx } = 3y + 4x^3\] \[\frac{ dy }{ dx } = \frac{ 3y + 4x^3 }{ 8y^3-3x }\]

OpenStudy (anonymous):

Check your signs

OpenStudy (wiggler):

\[-4x^3-3y = 3x \frac{ dy }{ dx } -8y^3\frac{ dy }{ dx }\]

OpenStudy (wiggler):

Then -4x^3 - 3y / 3x - 8y^3 = dy/dx.

OpenStudy (wiggler):

Which is right. Why does it only work moving the dx/dy to the right first? Or does it not matter and the other answer would be correct anyway?

OpenStudy (anonymous):

it's -3xdy/dx not -3y

OpenStudy (wiggler):

Where?

OpenStudy (anonymous):

You wrote -4x^3-3y

OpenStudy (anonymous):

Oh hold on I see what you're doing

OpenStudy (anonymous):

\[\huge -4x^3-3x \frac{ dy }{ dx }+8y^3\frac{ dy }{ dx }=3y\]

OpenStudy (anonymous):

You're just mixing up the algebra

OpenStudy (wiggler):

How would you do it?

OpenStudy (anonymous):

\[\huge \huge -4x^3+8y^3\frac{ dy }{ dx }=3y+3x \frac{ dy }{ dx }\] Ok you got to this point, agree?

OpenStudy (wiggler):

yep.

OpenStudy (anonymous):

Ok so isolate dy/dx by moving it to the left side.

OpenStudy (wiggler):

And the -4x^3 to the right?

OpenStudy (anonymous):

Yes, but with these problems, take it step by step, it's very easy to make mistakes.

OpenStudy (wiggler):

Okay :)

OpenStudy (wiggler):

Okay :)\[-4x^3 + 8y^3\frac{ dy }{ dx } - 3x \frac{ dy }{ dx } = 3y\]

OpenStudy (anonymous):

Good, continue :)

OpenStudy (wiggler):

\[8y^3\frac{ dy }{ dx } - 3x \frac{ dy }{ dx } = 3y + 4x^3\]

OpenStudy (wiggler):

Then isolating dy/dx:

OpenStudy (anonymous):

Correct

OpenStudy (wiggler):

\[(8y^3-3x)\frac{ dy }{ dx } = 3y + 4x^3\]

OpenStudy (anonymous):

Good! Finish it off!

OpenStudy (wiggler):

\[\frac{ dy }{ dx } = \frac{ 3y + 4x^3 }{ 8y^3 - 3x }\]

OpenStudy (anonymous):

Good job!

OpenStudy (wiggler):

Thank you!

OpenStudy (anonymous):

Your welcome :)!